# LangTraj: Diffusion Model and Dataset for Language-Conditioned Trajectory Simulation

### v1

@@ A.2 ProSim-Instruct-520k
A.3 Testing Subsets
@@
A.4 WOSAC Challenge Metrics
@@
$M_{K}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}w_{j}m_{K}(i,j),\quad\sum_{j=1}^{M}w_{j}=1$ (6)
where $M=9$ represents the nine statistical features, and $w_{j}$ are manually assigned weights for each metric. Safety-critical metrics, such as collisions and road departures, receive higher weights to emphasize their importance.
@@
Dataset Text Meta $\uparrow$ Kinematic $\uparrow$ Interactive $\uparrow$ Map $\uparrow$ mADE $\downarrow$
InterDrive None 0.72 0.41 0.80 0.80 2.65
Human-Annotated 0.72 0.41 0.80 0.80 2.73
Heuristic 0.72 0.42 0.80 0.79 2.36
Heuristic + Human-Annotated 0.72 0.42 0.80 0.79 2.29
Heuristic + Human-Annotated (interactive only) 0.72 0.42 0.80 0.79 2.31
Table 7: Evaluation of InterDrive. We analyze different levels of text conditioning, ranging from no text to heuristic and human-annotated labels. Combining heuristic and human annotations achieves the best performance in terms of minADE while maintaining strong interactive and map adherence metrics.
@@ B.2 Closed-loop Training Details
We modify the model’s prediction horizon from $T=16$ to $T=8$ to accommodate multi-step unrolling with ground truth actions, using a replanning interval of $T_{\text{replan}}=2$. Given 16 steps of ground truth future trajectories, we can perform four iterations of closed-loop training. During this process, the best sample among $M=8$ candidates is selected for execution, and the adopted forward diffusion ratio $\gamma=0.6$.
The teacher-forcing ratio is set to 50%, where 70% of agents are randomly sampled to follow the ground truth states throughout the unrolling process. The model is trained with a learning rate of $1\times 10^{-5}$ using an effective batch size of 32 for around 50,000 iterations.
Additionally, we incorporate an auxiliary non-collision loss from Sec. C.3 with a relative weighting of 0.1. To enhance robustness, we randomly drop text conditioning and agent history with a probability of 50% during training.
B.3 Denoising Process
@@
B.4 Diffusion Process and Inference Details
For the diffusion process, we utilize a cosine variance schedule, with the number of diffusion steps set to $K=100$ for pretrained diffusion model and $K=5$ for the closed-loop trained diffusion models. The cosine variance scheduler followed [19], with $s=0.008$. The model operates on a 1.1-second trajectory history and is trained to predict the next 8.0 seconds with a time step $\Delta t=0.5$. During inference, we use a DDIM sampler [27] with a stride step 1 during inference. During inference, we sample $M=64$ joint future samples for all agent, and only select the one joint agent sample with lowest collision loss.
Appendix C Guidance Details
C.1 Classifier-Free Guidance
@@
As shown in Tab. 6, we observe that moderate classifier-free guidance (weights 0.0–1.0) maintains realism and controllability, while higher weights (¿1.0) degrade map adherence (0.81 → 0.73) and interactive realism (0.79 → 0.76), suggesting that excessive text conditioning misaligns trajectories with the map structure.
C.2 Collision Guidance
@@ C.3 Non-Collision Guidance
Appendix D InterDrive Interaction Type Definitions

### v2

@@ A.1 InterDrive
During training, we also use compositional prompts that combine agent-agent interaction descriptions with heuristic action labels (e.g., speed up, turn left, wait) into unified instructions. This compositionality supports the flexibility of behavior expression, and we apply it consistently across both training and evaluation. Note that our instructions do not include explicit temporal conditioning, which we leave for future work.
@@ A.2 ProSim-Instruct-520k
A.3 LLM-Based Guidance Details
In this section, we provide implementation details for LLM-based guidance conditioning (CTG++ style [34]) method in Tab. 4, which generates differentiable loss functions conditioned on text descriptions. We use the o3-mini model via OpenAI APIs to generate loss functions that guide vehicle behaviors, following the implementation. The method leverages the same backbone and weights for all experiments to ensure consistency.
Since InterDrive uses a fixed vocabulary, we generate a unique function for each interactive description and heuristic, and combine them by scaling the loss to a common range. LLM-based guidance may not work in the first iteration, as the generated functions often contain errors or inconsistencies. Common failure cases include issues with array shape mismatches, map-related functions, and the assumption of unseen functions. To address this, we manually correct the generated functions by providing more specific instructions to GPT. This process typically requires 3-5 cycles to refine the guidance functions and ensure there are no compilation errors.
A.4 Testing Subsets
@@
A.5 WOSAC Challenge Metrics
@@
$M_{K}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}w_{j}m_{K}(i,j),\qquad\sum_{j=1}^{M}w_{j}=1$ (6)
where $M=9$ represents the nine statistical features, and $w_{j}$ are manually assigned weights for each metric. We detail the definitions of the component metrics below:
Kinematic Metrics:
  • Linear Speed: Measures the magnitude of the first derivative of position, $\|v\|=\left\|\frac{x_{t+1}-x_{t}}{\Delta t}\right\|_{2}$, reflecting the agent’s velocity in 3D space.
  • Linear Acceleration Magnitude: Represents the magnitude of the second derivative of position, $\frac{\|v_{t+1}-v_{t}\|}{\Delta t}$, describing the agent’s acceleration.
  • Angular Speed: Calculates the rate of change of the agent’s heading, $\omega=\frac{d(\theta_{t+1},\theta_{t})}{\Delta t}$, where $d(\cdot)$ is the minimal angular difference on the unit circle.
  • Angular Acceleration Magnitude: Measures the rate of change of angular speed, $\frac{d(\omega_{t+1},\omega_{t})}{\Delta t}$.
Interaction Metrics:
  • Distance to Nearest Object: The signed distance to the nearest object in the scene, calculated using the GJK distance algorithm.
  • Collisions: Detected when the signed distance to the nearest object becomes negative, indicating that two objects have collided.
  • Time-to-Collision (TTC): Estimates the time before a collision occurs, assuming constant velocities.
Map Metric:
  • Distance to Road Edge: The signed distance to the nearest road edge in the scene.
  • Road Departure: Indicates whether an agent has left the road at any point in time, based on the signed distance to the road edge.
For more details on each metric, refer to the original WOSAC Challenge paper [14].
@@ B.2 Closed-loop Training Details
We modify the model’s prediction horizon from $T=16$ to $T=8$ under 2 Hz to accommodate multi-step unrolling with ground truth actions, using a replanning interval of $T_{\text{replan}}=2$. Given 16 steps of ground truth future trajectories, we can perform four iterations of closed-loop training. During this process, the best sample among $M=8$ candidates is selected for execution, and the adopted forward diffusion ratio $\gamma=0.6$.
The teacher-forcing ratio is set to 50%. When applied, 70% of agents are randomly sampled to follow the ground truth states throughout the unrolling process. The model is trained with a learning rate of $1\times 10^{-5}$ using an effective batch size of 32 for around 50,000 iterations, which takes around 12 hours on 8 8×A6000 GPUs and 32 CPU cores.
Additionally, we incorporate an auxiliary non-collision loss from Sec. D.3 with a relative weighting of 0.1. To enhance robustness, we randomly drop text conditioning and agent history with a probability of 50% during training.
B.3 Inference Frequency
Per-sample inference with 5 de-noising steps takes 4.66±0.06 seconds per 8-second simulation (1 Hz replan) using 1×A6000 GPU and 4 CPU cores.Speedups via map caching, parallelism, and distillation can 069 be adopted for scalability.
B.4 Denoising Process
@@
B.5 Diffusion Process and Inference Details
For the diffusion process, we utilize a cosine variance schedule, with the number of diffusion steps set to $K=100$ for pretrained diffusion model and $K=5$ for the closed-loop trained diffusion models. The cosine variance scheduler followed [15], with $s=0.008$. The model operates on a 1.1-second trajectory history and is trained to predict the next 8.0 seconds with a time step $\Delta t=0.5$. During inference, we use a DDIM sampler [23] with a stride step 1 during inference. During inference, we sample $M=64$ joint future samples for all agent, and only select the one joint agent sample with lowest collision loss. The per-sample inference time is 4.66±0.06 seconds for an 8-second simulation, with a 1 Hz replan rate. This process utilizes a 1×A6000 GPU and 4 CPU cores. To improve scalability, speedups can be achieved through techniques such as map caching, parallelism, and distillation.
Appendix C Discussion and Limitation
While the model generally follows instructions well, we observe that failure cases often involve conflicts between language input and recent history (e.g., past 1s of motion). In such cases, the model tends to prioritize history, leading to behavior misaligned with the instruction. This is particularly noticeable when the vehicle is static, making conditioning signals harder to take effect. Additionally, minADE may not fully capture instruction adherence; future work could explore human evaluations to better assess language-action alignment.
Appendix D Guidance Details
D.1 Classifier-Free Guidance
@@
As shown in Tab. 7, we observe that moderate classifier-free guidance (weights 0.0–1.0) maintains realism and controllability, while higher weights (1.0) degrade map adherence (0.81 → 0.73) and interactive realism (0.79 → 0.76), suggesting that excessive text conditioning misaligns trajectories with the map structure. Qualitatively, while stronger guidance improves instruction-following, it also tends to produce more off-road samples.
D.2 Collision Guidance
@@ D.3 Non-Collision Guidance
Appendix E InterDrive Interaction Type Definitions